A positive interger ,n, is picked at random. Find the probability that the GCF of n and 60 is 2. Find the probability that that the GCF of n and 60 is not 1.
Found 2 solutions by robertb, richard1234:
Answer by robertb(5830)

So n is not from 1 to 500 inclusive now?
The problem is ill-posed. You have to limit the population set to a finite set, just like what you did in your first problem post 378532, from 1 to 500. Let me demonstrate: If n is any positive integer of the form n = 2p, where p is a prime not equal to 2,3,or 5, then exactly GCF(60,n) = 2. The probability would then be equal to (number of all positive primes not equal to 2,3,5)/(number of all positive integers). This ratio will not have an answer, and we have already restricted ourselves to the domain of the positive integers!

Answer by richard1234(7193)

For the GCF to equal 2, n must divide 2, and n must not divide 3 nor 5. First, note that there is a 1/2 probability of n being even. Now we will consider the set of even integers 2, 4, 6, 8, ... and compute the probability that a chosen number *will* be divisible by 3 or 5 (then subtract from 1).
By inclusion-exclusion principle, the probability that a number n is divisible by 3 or 5 is 1/3 + 1/5 - 1/15, or 7/15. 1 - 7/15 = 8/15, times 1/2 gives us a probability of 4/15.

For the GCF of n and 60 to not equal 1, n must divide any of 2, 3, or 5. This requires some inclusion-exclusion bashing.
We can compute this by P(2|n) + P(3|n) + P(5|n) - P(2,3|n) - P(2,5|n) - P(3,5|n) + P(2,3,5|n)
= 1/2 + 1/3 + 1/5 - 1/6 - 1/10 - 1/15 + 1/60
= 43/60
I hope that's the right answer :)